OpenStudy (anonymous):

log 12 ( v 2 + 35) = log 12 ( −12v − 1)

OpenStudy (shamil98):

\[\huge \log_{12} (v^2 +35) = \log_{12} (-12v-1)\] you can simplify this to:\[\huge v^2 + 35 = -12v -1\] \[\huge v^2 +12v + 36 = 0\] \[\huge (x+6)(x+6) = 0 , x = -6\]

OpenStudy (anonymous):